OpenStudy (anonymous):

How do you write this system of equations in matrix form? x²+y²=17 x+y=5

OpenStudy (anonymous):

I tried this, but that can't be right: [ 1 1 | 17 ] [ 1 1 | 5 ]

OpenStudy (owlcoffee):

I might be wrong on this one but: \[x^2+y^2=17\] \[x+y=5\] Then: \[x.x+y.y=17\] \[x+y=5\] \[\iff \left[\begin{matrix}x & y & 17 \\ 1 & 1 & 5\end{matrix}\right]\]

OpenStudy (owlcoffee):

Are we speaking about Cramers method of reduction?

OpenStudy (anonymous):

Hmm, no, not Cramers. Gauss-Jordan's elimination actually.

OpenStudy (trojanpoem):

Gauss elimination works with linear equations and the second one isn't , so somehow we have to turn it into one

OpenStudy (trojanpoem):

x²+y²=17 <<< if it was x^2 - y^2 It would have been easier.

OpenStudy (anonymous):

Certain. I solved it using regular algebra, the results were 1 and 4. Was just wondering if it's possible using linear algebra too.

OpenStudy (trojanpoem):

HOw about this ? (x+y) ^2 = 25 x^2 + y^2 + 2xy = 25 let x^2 = t , y^2 = k , xy = c t + k + 2c = 25 -> first eqn x^2+y^2 = 17 t + k = 17 -> 2nd

OpenStudy (trojanpoem):

Now we can put it into a matrix.

OpenStudy (trojanpoem):